# Geodesic Slice Sampler for Multimodal Distributions with Strong Curvature

### v1

@@ 5 Experiments
We evaluate MAGSS for targets with sharp curvature (Section 5.1), multiple modes (Section 5.2), or both (Section 5.3), always considering different choices of the metric. We also empirically quantify the effect of the numerical integrator 111 Code at: github.com/williwilliams3/magss .
@@
We use targets with known reference samples, which allows measuring the accuracy using the 1-Wasserstein (earth mover’s) distance with the samples provided by the algorithm [Flamary et al., 2021]. Besides accuracy, we quantify the samplers with the probability of jumping between the different modes, as the raw ratio of consecutive samples that are within separate modes (defined manually for each problem).
@@
To showcase the effect of the metric we will be running MAGSS also with in Euclidean metric, with $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})=\boldsymbol{I}_{D}$, and we additionally compared against parallel tempering and diffusive Gibbs sampling.
Parallel Tempering (PT) runs many Markov Chains in parallel, each of which has $p(\operatorname{\boldsymbol{x}})^{1/\tau_{i}}$ as targets for different temperatures $\tau_{i}\geq 1$, where for $\tau_{1}=1$ the original target is recovered. As $\tau\to\infty$, the density flattens, facilitating transitions between regions of higher densities that are far apart from each other. The parallel chains jumps randomly between each other, thus visiting the modes more often according to a Metropolis-Hastings ratio [Swendsen and Wang, 1986, Geyer, 1991]. Our implementation of PT follows Łatuszyński et al. [2025].
Diffusive Gibbs sampling (DiGS) by Chen et al. [2024] is a sampler designed for addressing multi-modality. It approaches the sampling task by using an auxiliary variable with a Gibbs scheme. It uses the variance preserving (VP) [Song et al., 2021] noise scaling: $p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})=\operatorname{\mathcal{N}}(\tilde{\operatorname{\boldsymbol{x}}}|\alpha_{t}\operatorname{\boldsymbol{x}},\sigma^{2}_{t})$, where $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$, sampled directly and $p(\operatorname{\boldsymbol{x}}|\tilde{\operatorname{\boldsymbol{x}}})\propto p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})p(\operatorname{\boldsymbol{x}})$ sampled through a local MCMC sampler. It has an additional Metropolis within Gibbs proposal scheme $q(\operatorname{\boldsymbol{x}}|\tilde{\operatorname{\boldsymbol{x}}})=\operatorname{\mathcal{N}}(\operatorname{\boldsymbol{x}}|\tilde{\operatorname{\boldsymbol{x}}}/\alpha_{t},(\alpha_{t}/\sigma_{t})^{2})$. VP has the property that at when $\alpha_{t}\to 0$ then $p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})=\operatorname{\mathcal{N}}(\tilde{\operatorname{\boldsymbol{x}}}|0,\operatorname{\boldsymbol{I}}_{D})$ and when $\alpha_{t}\to 1$ then $p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})=\delta_{\operatorname{\boldsymbol{x}}}$.
@@ 5.1 Complex unimodal targets
We evaluate the methods on three canonical benchmark targets (funnel, hybrid Rosenbrock and squiggle) with strong curvature, with the densities given in Appendix C.5. PT is left out here and we only consider the metrics in Section 4.1 since the target is unimodal.
@@
We obtain $10,000$ samples using $10$ chains and omit results for runs that did not complete in 12 hours. We set $\alpha^{2}=1$ for the inverse Monge metric since this value has been shown to work in 2 dimensions [Hartmann et al., 2022]. We select $\lambda=1$, $p_{0}=1$ and inverse Generative without further tuning. We use Dopri5 integrator with adaptive step-size. We set $w=3$ and $m=8$. DiGS uses a single noise scale $\alpha=1$ and step-size $0.1$ for MALA within the algorithm.
@@ 5.2 Multimodal with simple modes
For studying mode exploration, we use a target of two $D$-dimensional Gaussian distributions centered at $-\boldsymbol{1}_{D}$ and $\boldsymbol{1}_{D}$ with scale $\sigma=0.1$ and weights $\{0.2,0.8\}$. The distance between the modes ($\sqrt{D}2$) grows for higher dimensions, making transition between the modes more difficult. Now we only consider the new metrics for boosting mixing between the modes (Section 4.2).
Table 1 reports the percentage of jumps between the modes and Figure 5 reports the corresponding accuracies. While the comparison methods PT and DiGS explore the modes well in low dimensions, they get completely stuck in one mode for $D\geq 8$. MAGSS with the inverse Monge metric ($\alpha=0.1$) is able to jump between the modes even for higher dimensions and is overall the most accurate method. The meta-sampler retains efficient exploration throughout the experiment, but fails to reach the correct balance between the two modes, which is shown as higher Wasserstein distance.
Table 1: Mixture of two Gaussians ($\operatorname{\boldsymbol{G}}_{IM}$, with $\alpha=0.1$).
jumpdimension
sampler 2 4 8 16 32 64
MAGSS 8.96 5.07 2.28 0.8 0.2 0.03
Meta-MAGSS 18.91 12.33 7.5 4.29 2.45 1.07
DiGS 5.05 0.02 0.0 0.0 0.0 0.0
PT 12.54 4.68 0.23 0.0 0.0 0.0
@@ 5.3 Multimodal with complex modes
To demonstrate that we can simultaneously handle multimodality and complex local geometry, we consider a (uniform) mixture of two narrow bivariate distributions, the Rosenbrock and Squiggle distributions (Figure 6 left; the red line is purely for identifying jumps between the modes). We use use the inverse Monge and inverse Generative metrics. However, Figure 6 (right) indicates that PT is the least accurate method, requiring substantially more samples for matching the target well. All methods will reach approximately the same Wasserstein distance if ran long enough, but both of our variants achieve it in less samples, confirming more efficient mixing.
@@
MAGSS $\operatorname{\boldsymbol{G}}_{Ig}$, $\lambda=1.0$ $0.82$ 17
Meta-MAGSS $\operatorname{\boldsymbol{G}}_{IM}$, $\alpha^{2}=10^{-4}$ $0.99$ 134
@@
Obtain $10,000$ samples using $10$ chains. We run DiGS with 5 noise steps between $0.1$ and $0.9$, and $10$ MALA iterations per sample. PT uses $\tau\in\{1.0,5.62,31.62,177.83,1000\}$ and thinning of $10$. MAGSS and Meta-MAGSS  are tuned using the same grid of values as Experiment 5.2, reporting the best based on distances. Meta-MAGSS uses $5$ sweeps, $10$ MALA iterations per sample. PT, DiGS and Meta-MAGSS rely on MALA with stepsize $0.001$ (roughly 60% acceptance rate). We use $w=3$ and $m=8$ and the Dopri8 integrator with adaptive step-size.
Figure 6: Left: Mixture of narrow distributions, with samples using Meta-MAGSS. Right: Wasserstein distance as a function of iterations (samples). For MAGSS we use $\operatorname{\boldsymbol{G}}_{IM}$ with $\alpha^{2}=10^{-4}$ and for Meta-MAGSS $\operatorname{\boldsymbol{G}}_{Ig}$ with $\lambda=1$.
@@ 5.4 Effect of numerical integrator
We use numerical integrators for computing the geodesics in Eq. (2). To explore the effect of the integrator, we present results for broad range of integrators for a multimodal benchmark task considered previously by Chen et al. [2024]: The target is a 40-mode Gaussian mixture model with equal weights and each component of variance $\sigma=0.1$ where the means are distributed uniformly on the square $(-40,40)^{2}$.
We use seven different integrators for $\operatorname{\boldsymbol{G}}_{IM}$ and $\operatorname{\boldsymbol{G}}_{Ig}$ metrics, including both adaptive and fixed step-sizes as implemented in [Kidger, 2021], and report the results in Figure 7. The three main conclusion are: (a) For metrics that are further away from Euclidean (large $\alpha$ or small $\lambda$) the integration time for adaptive methods grows dramatically. This is a natural consequence of operating in a less flat geometry. (b) For good accuracy we typically need to use such a geometry, which means there is inherent compromise between accuracy an computation. (c) Simple fixed-step integrators, even the Euler method, are efficient when they work, but for robustness we recommended adaptive methods. We recognize dopri5 as a good practical recommendation, but Euler is worth trying for the Inverse Generative metric.
@@ 6 Related work
Lan et al. [2014] constructed the Wormhole Hamiltonian Monte Carlo where a specific geometry is constructed to (only) connect the modes of multi-modal distributions. The modes are first identified along the Markov Chain evolution. After a new mode identification, it "stores" the mode’s location for later use. A jump using the updated mode candidates guarantees correct detailed balance equations. While this work served as an inspirational motivation for us, it requires notable additional components. MAGSS does not require separate identification or storage of the modes, but instead shrinks the distances by naturally warping the space.
@@ 7 Conclusions
Our aim was to show that multimodality and local curvature can be address by the same set of tools, namely Riemannian geometry. We provided a concrete Riemannian slice sampler, introduced two new metrics for improving mixing between modes, and showed that we can achieve accuracy and mixing comparable to recent samplers designed specifically for multi-modal targets, by only using a metric for this.
The obvious limitation is the computational cost. We form geodesics by numeric integration and even when using metrics with closed-form inverses the per-iteration cost of MAGSS is larger than of competing methods. However, we note that we used maximally exact solvers rather than seeking for the highest computational efficiency. Now that the principle has been demonstrated, use of more approximative numerical integrators for speeding up the overall computation could be studied in future work.
@@ Acknowledgements.
The work was supported by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence (FCAI), and additionally by grants: 363317, 345811 and 348952. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.
@@ A.1 Meta-Metric-agnostic Geodesic Slice Sampler
Algorithm 2 is the combination of MAGSS for $K$-steps followed by a local MCMC sampler for $L$-steps.
@@
4:         Update $\operatorname{\boldsymbol{x}}$ by Approximate Geodesic Slice Sampling starting at $\operatorname{\boldsymbol{x}}$
@@
7:         Update $\operatorname{\boldsymbol{x}}$ by local MCMC sampler starting at $\operatorname{\boldsymbol{x}}$.
@@ A.2 Step-out and Shrinkage procedures
The stepping-out and shrinkage procedures can be found in algorithm boxes 3 and 4, the algorithms are taken from Durmus et al. [2023]. The code implementation of the step-out procedure has vectorized both while loops by evaluation of the log density on all step-out points at once. The code implementation of the shrinkage procedure has a max number of iteration set at $100$ of the while loop, which defaults back to the previous point of the chain if exceeded. In the algorithm boxes we use the notation for the exponential map $\gamma_{(x,v)}(t)$.

### v2

@@ 5 Experiments
We evaluate MAGSS for targets with sharp curvature (Section 5.1), multiple modes (Section 5.2), or both (Section 5.3), always considering different choices of the metric. We also empirically quantify the effect of the numerical integrator. A code reproducing the experiments is available at github.com/williwilliams3/magss.
@@
We use primarily targets with known reference samples, which allows measuring the accuracy using the 1-Wasserstein (earth mover’s) distance with the samples provided by the algorithm [Flamary et al., 2021]. Besides accuracy, we quantify the samplers with the probability of jumping between the different modes, as the raw ratio of consecutive samples that are within separate modes (defined manually for each problem).
@@
To showcase the effect of the metric we will be running MAGSS also with in Euclidean metric, with $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})=\boldsymbol{I}_{D}$, and we additionally compared against the No-U-Turn Sampler, parallel tempering and diffusive Gibbs sampling.
The No-U-Turn Sampler (NUTS) is an auxiliary-variable sampler that augments the position $\operatorname{\boldsymbol{x}}_{t}$ with a velocity $\operatorname{\boldsymbol{v}}_{t}$ which jointly follow the Hamiltonian dynamics [Neal et al., 2011]. It adaptively determines the integration time by stopping at the first U-turn, i.e., the first time $t>0$ such that $\langle\operatorname{\boldsymbol{x}}_{t}-\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{t}\rangle<0$ [Hoffman et al., 2014].
Parallel tempering (PT) runs many Markov Chains in parallel, each of which has $p(\operatorname{\boldsymbol{x}})^{1/\tau_{i}}$ as targets for different temperatures $\tau_{i}\geq 1$, with $\tau_{1}=1$ recovering the original target. As $\tau\to\infty$, the density flattens, facilitating transitions between regions of higher densities that are far apart from each other. The parallel chains jumps randomly between each other, thus visiting the modes more often according to a Metropolis-Hastings ratio [Swendsen and Wang, 1986, Geyer, 1991]. Our implementation of PT follows Łatuszyński et al. [2025].
Diffusive Gibbs sampling (DiGS) by Chen et al. [2024] is a sampler designed for addressing multimodality. It approaches the sampling task by using an auxiliary variable $\tilde{\operatorname{\boldsymbol{x}}}$ with a Gibbs scheme. It uses the variance preserving (VP) [Song et al., 2021] noise scaling: $p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})=\operatorname{\mathcal{N}}(\tilde{\operatorname{\boldsymbol{x}}}|\alpha_{t}\operatorname{\boldsymbol{x}},\sigma^{2}_{t})$, where $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$, sampled directly and $p(\operatorname{\boldsymbol{x}}|\tilde{\operatorname{\boldsymbol{x}}})\propto p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})p(\operatorname{\boldsymbol{x}})$ sampled through a local MCMC sampler. It has an additional Metropolis within Gibbs proposal scheme $q(\operatorname{\boldsymbol{x}}|\tilde{\operatorname{\boldsymbol{x}}})=\operatorname{\mathcal{N}}(\operatorname{\boldsymbol{x}}|\tilde{\operatorname{\boldsymbol{x}}}/\alpha_{t},(\alpha_{t}/\sigma_{t})^{2})$. VP has the property that at when $\alpha_{t}\to 0$ then $p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})=\operatorname{\mathcal{N}}(\tilde{\operatorname{\boldsymbol{x}}}|0,\operatorname{\boldsymbol{I}}_{D})$ and when $\alpha_{t}\to 1$ then, informally, $p(\tilde{\operatorname{\boldsymbol{x}}}|\operatorname{\boldsymbol{x}})=\delta_{\operatorname{\boldsymbol{x}}}$.
@@ 5.1 Complex unimodal targets
We evaluate the methods on three canonical benchmark targets (funnel, hybrid Rosenbrock and squiggle) which exhibit strong curvature. The densities are given in Appendix C.5. Since these targets are unimodal, we only consider the metrics presented in Section 4.1 and exclude PT.
@@
We obtain $10,000$ samples using $10$ chains and omit results for runs that did not complete in 12 hours. We set $\alpha^{2}=1$ for the Monge metric since this value has been shown to work [Hartmann et al., 2022]. We select $\lambda=1$, $p_{0}=1$ for the Generative metric without further tuning. We use Dopri5 integrator with adaptive step-size. We set $w=3$ and $m=8$. DiGS and NUTS uses a single noise scale $\alpha=1$ and step-size $0.1$ for MALA within the algorithm.
@@ 5.2 Multimodal with simple modes
For studying mode exploration, we use a target of two $D$-dimensional Gaussian distributions centered at $-\boldsymbol{1}_{D}$ and $\boldsymbol{1}_{D}$ with scale $\sigma=0.1$ and weights $\{0.2,0.8\}$. The distance between the modes ($\sqrt{D}2$) grows for increasing dimensions, making transition between the modes more difficult. Now we only consider the new metrics for boosting mixing between the modes (Section 4.2).
Figure 5 reports the corresponding accuracies and reports the percentage of jumps between the modes. While the comparison methods PT and DiGS explore the modes well in low dimensions, they get completely stuck in one mode for $D\geq 16$. MAGSS and Meta-MAGSS with the Inverse Monge metric ($\alpha=0.1$) are able to jump between the modes even for higher dimensions and the meta-sampler is overall the most accurate method.
@@ 5.3 Multimodal with complex modes
To demonstrate that we can simultaneously handle multimodality and complex local geometry, we consider a (uniform) mixture of two narrow bivariate distributions, the Rosenbrock and Squiggle distributions (Figure 6 left; the red line is purely for identifying jumps between the modes, Table 1). We use use the Inverse Monge and Inverse Generative metrics. However, Figure 6 (right) indicates that PT is the least accurate method, requiring substantially more samples for matching the target well. All methods will reach approximately the same Wasserstein distance if ran long enough, but both of our variants achieve it in less samples, confirming more efficient mixing.
@@
MAGSS $\operatorname{\boldsymbol{G}}_{Ig}$, $\lambda=1.0$ $0.81$ 178
Meta-MAGSS $\operatorname{\boldsymbol{G}}_{IM}$, $\alpha^{2}=10^{-4}$ $2.62$ 1327
@@
Obtain $10,000$ samples using $10$ chains. We run DiGS with 5 noise steps between $0.1$ and $0.9$, and $10$ MALA iterations per sample. PT uses $\tau\in\{1.0,5.62,31.62,177.83,1000\}$ and thinning of $10$. MAGSS and Meta-MAGSS  are tuned using the same grid of values as Experiment 5.2, reporting the best based on distances. Meta-MAGSS uses $5$ sweeps and $10$ MALA iterations per sample. PT, DiGS and Meta-MAGSS rely on MALA with stepsize $0.001$ ($\approx 60\%$ acceptance rate). We use $w=3$ and $m=8$ and the adaptive Dopri8 integrator.
Figure 6: Left: Mixture of narrow distributions, with samples using Meta-MAGSS. Right: Wasserstein distance as a function of iterations (samples). For MAGSS we use $\operatorname{\boldsymbol{G}}_{Ig}$ with $\lambda=1$ and Meta-MAGSS $\operatorname{\boldsymbol{G}}_{IM}$ with $\alpha^{2}=10^{-4}$.
5.4 Field System
We include a highly multimodal target distribution of modes ($2^{D}$ for $D=16$) by replicating the Allen-Cahn Field System model experiment from Cabezas et al. [2024b]. The distribution has two global maxima at (1,..,1)(1,..,1)( 1 , . . , 1 ) and (1,..,1)(-1,..,-1)( - 1 , . . , - 1 ), and several lower density modes at points of the form $x_{i}=\pm 1$ for all $i$. DiGS collapses to a single mode, PT explores only the two most dominant modes, while our Meta-MAGSS explores also the additional modes (Fig 7).
The target density is symmetric along each axis. The initial sampling position is $(-1,\ldots,-1)$ and we use the marginal distribution of $x_{8}$ to evaluate how well each sampler captures the symmetry. In particular, we report the percentage of samples with $x_{8}>0$, which should be $50\%$ under the true distribution. Since reference samples are not directly available, we follow Cabezas et al. [2024b] and also report the Kernel Stein Discrepancy (KSD V-stat) [Liu et al., 2016] in Table 2.
Our method explores more modes than the competing methods (PT and DiGS), although the KSD V-stat is worse. Note, however, that the KSD V-stat does not account for multimodality at all; DiGS has a better value despite covering only a single mode and failing completely in terms of the marginal distribution metric. In contrast, PT and Meta-MAGSS exhibit similar percentages of samples with $x_{8}>0$. We provide the density of the model, an explanation of the multimodality of the model, and the computation of KSD V-stat in Appendix C.5.
Table 2: Field System model
sampler KSD V-stat $x_{8}>0$ t(s)
PT $0.13\pm 0.04$ $0.35$ 6
DiGS $0.13\pm 0.05$ $0.0$ 157
META-AGSS $2.98\pm 0.7$ $0.33$ 32
Experiment specifications:
We obtain $10,000$ samples using $10$ chains initialized at (1,..,1)(-1,..,-1)( - 1 , . . , - 1 ). DiGS uses $T=1000$ equally spaced times between $\alpha_{1}=10^{-5}$, $\alpha_{1000}=1-10^{-5}$. PT uses $N=200$ temperatures in the scale $\tau_{i}=b_{\min}^{-{i}/{N}}$ for $i=1,..,N$ where $b_{\mathrm{min}}=10^{-5}$. For Meta-MAGSS we try values of $\alpha$ and $\lambda$ in powers of ten, finding $\lambda=10^{-6}$ maximizes the number of jumps between modes and a single sweep. We use $w=3$ and $m=8$ and the Dopri5 integrator with adaptive step-size. For all methods MALA uses $10$ iterations per sample and stepsize $0.005$ (roughly 60% acceptance rate).
Figure 7: Samples from Allen-Cahn Field System model [Cabezas et al., 2024b] with $2^{16}$ modes that zig-zag between $-1$ and $1$ on the y-axis over the $D=16$ values at the x-axis. Euclidean DiGS (left) gets stuck in one mode, Parallel Tempering (middle) only explores two dominant modes with constant value over the x-axis, whereas Meta-MAGGS (right; $\boldsymbol{G}_{Ig}$ with $\lambda=10^{-6}$) explores also the modes that switch between the extremes.
@@ 5.5 Effect of numerical integrator
We use numerical integrators for computing the geodesics in Eq. (2). To explore the effect of the integrator, we present results for broad range of integrators for a multimodal benchmark task considered previously by Chen et al. [2024]. The target is a 40-mode Gaussian mixture model with equal weights and each component of variance $\sigma=0.1$ where the means are distributed uniformly on the square $(-40,40)^{2}$.
We use seven different integrators for $\operatorname{\boldsymbol{G}}_{IM}$ and $\operatorname{\boldsymbol{G}}_{Ig}$ metrics, including both adaptive and fixed step-sizes as implemented by Kidger [2021], and report the results in Figure 8. The three main conclusion are: (a) For metrics that are further away from Euclidean (large $\alpha$ or small $\lambda$) the integration time for adaptive methods grows dramatically. This is a natural consequence of operating in a less flat geometry. (b) For good accuracy we typically need to use such a geometry, which means there is inherent compromise between accuracy an computation. (c) Simple fixed-step integrators, even the Euler method, are efficient when they work, but for robustness we recommended adaptive methods. We recognize dopri5 as a good practical recommendation, but Euler is worth trying for the Inverse Generative metric.
@@ 6 Related work
Lan et al. [2014] constructed the Wormhole Hamiltonian Monte Carlo where a specific geometry is built to (only) connect the modes of multimodal distributions. The modes are first identified along the Markov Chain evolution. After a new mode identification, it "stores" the mode’s location for later use. A jump using the updated mode candidates guarantees correct detailed-balance equations. While this work served as an inspirational motivation for us, it requires notable additional components. However, MAGSS does not require separate identification or storage of the modes, but instead shrinks the distances by naturally warping the space.
@@ 7 Conclusions
Our aim was to show that local curvature and multimodality can be addressed by the same set of tools, namely Riemannian geometry. We provided a concrete Riemannian slice sampler, introduced two new metrics for improving mixing between modes, and showed that we can achieve accuracy and mixing comparable to recent samplers designed specifically for multimodal targets, by only using Riemannian metrics for this task.
One obvious limitation is the computational cost, caused by numeric integration of the geodesics. Even when using metrics with fast inverses, the per-iteration cost of MAGSS is larger than of competing methods. However, we note that we used maximally exact solvers rather than seeking for the highest computational efficiency. Now that the principle has been demonstrated, the use of more approximative numerical integrators for speeding up the overall computation could be studied in future work.
@@ Acknowledgements.
BW, HY, HPHL and AK were supported by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence (FCAI), and additionally by grants: 363317, 345811 and 348952. GA was supported by the DFF Sapere Aude Starting Grant “GADL”. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.
Appendix
@@ A.1 Meta-Metric-agnostic Geodesic Slice Sampler
The Meta-MAGSSfound in algorithm 2 is the combination of MAGSS for $K$-steps followed by a local MCMC sampler for $L$-steps.
@@
4:         Update $\operatorname{\boldsymbol{x}}$ by MAGSS with initial position $\operatorname{\boldsymbol{x}}$
@@
7:         Update $\operatorname{\boldsymbol{x}}$ by local MCMC with initial position at $\operatorname{\boldsymbol{x}}$.
@@ A.2 Step-out and Shrinkage procedures
The stepping-out and shrinkage procedures are Algorithm 3 and Algorithm 4 respectively, these algorithms are taken from Durmus et al. [2023]. Our code implementation of the step-out procedure has vectorized both while loops in Algorithm 3. This is done by evaluating the log density on all possible step-out points at once (vectorized). The code implementation of the shrinkage procedure (Algorithm 4) has a max number of iteration set at $100$ for the while loop, which if exceeded defaults back to the previous point of the chain. In the algorithm boxes we use the notation for the exponential map $\gamma_{(x,v)}(t)$. JAX is used to handle automatic differentiation and the samplers are coded in the style of Blackjax [Bradbury et al., 2018, Cabezas et al., 2024a].